# Breaking Data Silos: Towards Open and Scalable Mobility Foundation Models via Generative Continual Learning

### v5

@@ 5.5. Ablation Studies
As shown in Figure 6, removing any input feature from the Mobility-Aware Expert Routing module leads to a noticeable performance drop. Similarly, disabling the knowledge distillation strategy in GCL also results in a significant decline in model performance. These findings highlight the critical role of each input feature in expert selection, as well as the importance of knowledge distillation in ensuring model stability during continual learning.
@@ 5.7. Impact of Replay Volume
As shown in Figure 8, the performance on the base cities (WSC) consistently improves with more generated data, indicating that replay volume directly affects the ability to retain past knowledge. In contrast, performance on the newly introduced cities (A, L, N) remains largely stable regardless of replay volume, with no consistent trend of improvement or degradation. These results suggest that while synthetic replay is crucial for mitigating forgetting, it has limited effect on new knowledge acquisition. Thus, allocating a moderate amount of generated data offers a practical balance—sufficient to preserve prior knowledge without incurring unnecessary overhead—supporting the long-term scalability.
@@ 6. Conclusion
In this work, we present MoveGCL, a scalable and privacy-preserving framework for training mobility foundation models via generative continual learning. By enabling decentralized model evolution without sharing raw data, it addresses key challenges in real-world human mobility modeling, including data silos, privacy constraints, and heterogeneous mobility distributions. MoveGCL represents a significant step toward realizing mobility foundation models by offering a practical and generalizable framework that facilitates collaborative learning across cities and institutions. It paves the way for long-term, privacy-safe, and adaptive modeling of human movement, with broad implications for urban planning, transportation optimization, and evidence-based policy making. The development of more large-scale, semantically rich, and geographically diverse mobility datasets will be crucial for further improving the model’s generalization and robustness. We encourage the broader research community and data-holding institutions to join this collaborative effort, contributing to the creation of open, inclusive, and powerful spatiotemporal foundation models for the mobility domain.
Acknowledgments
This work was supported in part by the National Key Research and Development Program of China under grant 2024YFC3307603 and the National Natural Science Foundation of China under 62476152.
Appendix A Dataset
@@
Appendix B Evaluation Metrics

### v6

@@ 5.5. Ablation Studies
As shown in Appendix Figure 7, removing any input feature from the Mobility-Aware Expert Routing module leads to a noticeable performance drop. Similarly, disabling the knowledge distillation strategy in GCL also results in a significant decline in model performance. These findings highlight the critical role of each input feature in expert selection, as well as the importance of knowledge distillation in ensuring model stability during continual learning.
@@ 5.7. Impact of Replay Volume
As shown in Appendix Figure 8, the performance on the base cities (WSC) consistently improves with more generated data, indicating that replay volume directly affects the ability to retain past knowledge. In contrast, performance on the newly introduced cities (A, L, N) remains largely stable regardless of replay volume, with no consistent trend of improvement or degradation. These results suggest that while synthetic replay is crucial for mitigating forgetting, it has limited effect on new knowledge acquisition. Thus, allocating a moderate amount of generated data offers a practical balance—sufficient to preserve prior knowledge without incurring unnecessary overhead—supporting the long-term scalability.
6. Limitations and Ethical Considerations
While MoveGCL provides a robust framework for bridging data silos in mobility science, it faces several limitations and ethical challenges. First, the scientific generalizability of the model is inherently tied to the diversity of participating institutions; if the collaborative training sequence is skewed toward megacities in developed regions, it may fail to accurately represent the unique mobility dynamics of the Global South or rural areas, potentially leading to inequitable urban planning recommendations. From an ethical standpoint, although generative replay provides a significant privacy buffer, it still necessitates future research into integrating formal Differential Privacy (DP) guarantees. Furthermore, high-fidelity mobility foundation models present dual-use concerns; while they are intended to advance sustainability goals such as carbon footprint reduction, strict governance is required to prevent their misuse for unauthorized surveillance or restrictive movement policies.
@@ 7. Conclusion
In this work, we present a scalable and privacy-preserving framework for training mobility foundation models via generative continual learning. By enabling decentralized model evolution without sharing raw data, it addresses key challenges in real-world human mobility modeling, including data silos, privacy constraints, and heterogeneous mobility distributions. MoveGCL represents a significant step toward realizing mobility foundation models by offering a practical and generalizable framework that facilitates collaborative learning across cities and institutions. It paves the way for long-term, privacy-safe, and adaptive modeling of human movement, with broad implications for urban planning, transportation optimization, and evidence-based policy making. The development of more large-scale, semantically rich, and geographically diverse mobility datasets will be crucial for further improving the model’s generalization and robustness. We encourage the broader research community and data-holding institutions to join this collaborative effort, contributing to the creation of open, inclusive, and powerful spatiotemporal foundation models for the mobility domain.
Appendix A Experimental Settings
A.1. Dataset
@@
A.2. Baselines
A.3. Parameter Settings
The key parameters of our framework fall into three main categories.
A.4. Evaluation Metrics
@@
Appendix B Related Work
Mobility Data
Mobility Foundation Models
Continual Learning
Continual learning (kim2022theoretical; wang2024comprehensive; wang2024comprehensive), also known as lifelong learning, aims to enable models to learn from a sequence of tasks or data streams without forgetting previously acquired knowledge (yoon2021federated). A central challenge in continual learning is catastrophic forgetting (li2019learn; wickramasinghe2023continual), where the model’s performance on earlier tasks degrades as it learns new ones. Continual learning methods are typically categorized into three main types. Regularization-based methods introduce constraints on parameter updates to preserve important knowledge from earlier tasks. Replay-based methods mitigate forgetting by either storing a subset of previous data (experience replay) or generating pseudo-data (generative replay) to simulate past learning. Parameter isolation methods allocate different parts of the model to different tasks, using techniques such as dynamic networks or task-specific masking to reduce interference between tasks.
Appendix C Additional Results
Model Atlanta Chicago Los Angeles New York Seattle Washington D.C
[…]
C.1. Order Invariance in Continual Learning
C.2. Experimental Settings of Privacy Evaluation
Uniqueness Testing
Membership Inference Attack
Differential Privacy
In our experiment, we randomly select a subset of trajectories and consider two training scenarios: one in which this subset is included in the training data ($D$), and one in which it is excluded ($D^{\prime}$). For each scenario, we train $M$ on the corresponding dataset and then use each selected trajectory as a conditioning input to generate multiple synthetic trajectories. We compute a similarity score between each generated trajectory and its original conditioning trajectory. These similarity scores are modeled as two Gaussian distributions corresponding to $\mathbb{P}\bigl[M(z;\,D)\bigr]$ and $\mathbb{P}\bigl[M(z;\,D^{\prime})\bigr]$, respectively. Finally, we estimate the privacy‐budget parameters $\varepsilon$ from these distributions.
C.3. Supplementary Figures